This question introduces a fundamental result of taxation which will revisit in the last chapter. We...

  1. This question introduces a fundamental result of taxation which will revisit in the last chapter. We can already see it at work through the following example:

A firm faces the following demand curve: P = 120 – 0.02Q

Where Q is weekly production and P is price, measured in cents per unit. The firm’s total cost function is given by TC = 60Q + 25,000. Assume that the firm maximizes profit.

  1. What is the level of production, price, and total profit per week?

(Hint: to answer the question, you need to know MR and MC. In this case, MR = 120 – 0.04Q and MC = 60. [These results can be found by way of calculus, but you are not responsible for knowing this.])

  1. If the government decides to levy a tax of 14 cents per unit on this product, what will be the new level of production, price, and profit?
Answer #1

2 P-120-o-oze MR-120 -0.oud an c=60 Prefit is mad when MR-me lao - 0.04 0=60 0=60 - 6000= 15004 oo4= - Pein = 120-0.02(1500)=